# Semantic Agreement Enables Efficient Open-Ended LLM Cascades

### v1

@@ H.2.4 CNN/DailyMail
Model Random Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile
Qwen2.5-7B 0.258 0.2577 0.2605 0.2607 0.2
Qwen2.5-3B 0.2511 0.2539 0.2537 0.2537 0.5
Qwen2.5-1.5B 0.2474 0.2407 0.2536 0.2559 1.0
Qwen2.5-0.5B 0.2418 0.2371 0.245 0.2467 1.0
Gemma3-1B 0.2448 0.2454 0.2473 0.2479 0.4
Llama3.2-1B 0.2482 0.2502 0.2508 0.2511 0.3
Llama3.2-3B 0.2522 0.2535 0.2548 0.2549 0.1
Llama3.1-8B 0.2551 0.2547 0.2564 0.2565 0.2
mT0-Large 0.2527 0.2528 0.2539 0.2542 0.8
@@
Ensemble Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile
Llama3.2-1B, mT0-Large 0.2554 0.2529 0.2553 0.1
Llama3.2-1B, Qwen2.5-1.5B 0.2532 0.2493 0.2548 0.2
Qwen2.5-1.5B, Gemma3-1B 0.2505 0.2489 0.2559 1.0
Llama3.2-1B, mT0-Large, Qwen2.5-1.5B 0.2568 0.2491 0.257 0.8
Qwen2.5-7B, Llama3.1-8B 0.2579 0.259 0.2611 0.2
Qwen2.5-7B, mT0-Large 0.258 0.2575 0.2588 0.2
Qwen2.5-7B, Qwen2.5-3B 0.259 0.2583 0.2602 0.2
@@
Ensemble BLEU ROUGE-1 ROUGE-2 ROUGE-L BLEURT SBERT
Llama3.2-1B, mT0-Large,
Qwen2.5-1.5B
0.2611 0.2616 0.2598 0.2608 0.2567 0.2603
Llama3.2-1B, mT0-Large,
Qwen2.5-0.5B
0.2566 0.2567 0.2553 0.2574 0.2513 0.2547
Qwen2.5-1.5B, Qwen2.5-0.5B,
Gemma3-1B
0.2551 0.2546 0.2514 0.2519 0.2522 0.2515
Llama3.2-1B, Gemma3-1B,
Qwen2.5-0.5B
0.2537 0.2538 0.2534 0.2543 0.251 0.2528
Llama3.2-3B, Qwen2.5-3B,
Qwen2.5-0.5B
0.2568 0.258 0.2578 0.2578 0.2546 0.2576
Qwen2.5-7B, mT0-Large,
Llama3.2-3B
0.2635 0.2618 0.2615 0.2607 0.2584 0.2618
Qwen2.5-7B, Llama3.1-8B,
mT0-Large
0.2622 0.2613 0.2622 0.2619 0.2586 0.2612
Llama3.1-8B, mT0-Large,
Qwen2.5-3B
0.2589 0.2571 0.2583 0.2578 0.2555 0.2585
@@
Model Random Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile
Qwen2.5-7B 0.8053 0.8217 0.8201 0.821 0.0
Qwen2.5-3B 0.7623 0.7916 0.7921 0.7915 0.1
Qwen2.5-1.5B 0.7593 0.7897 0.7847 0.7868 0.0
Qwen2.5-0.5B 0.642 0.6898 0.6826 0.685 0.0
Llama3.1-8B 0.7987 0.8173 0.8151 0.8169 0.0
Llama3.2-3B 0.7627 0.7764 0.7791 0.7833 0.5
Llama3.2-1B 0.631 0.6695 0.677 0.6768 0.3
mT0-Large 0.802 0.8205 0.8211 0.8211 0.1
FLAN-T5-Large 0.7123 0.7598 0.6896 0.7437 0.0
Gemma3-1B 0.676 0.6984 0.6938 0.696 0.0
@@
Ensemble Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile
mT0-Large, Qwen2.5-1.5B 0.8155 0.8168 0.8125 0.0
mT0-Large, Qwen2.5-1.5B, FLAN-T5-Large 0.7809 0.8081 0.8043 0.0
Qwen2.5-1.5B, FLAN-T5-Large, Gemma3-1B 0.7414 0.7511 0.7632 0.0
Qwen2.5-7B, Llama3.1-8B 0.8211 0.8211 0.8219 0.0
Qwen2.5-7B, Llama3.1-8B, mT0-Large 0.8236 0.8239 0.8255 0.0
Qwen2.5-7B, Llama3.2-3B 0.7993 0.7985 0.7988 0.0
Qwen2.5-7B, Llama3.2-3B, Gemma3-1B 0.7599 0.7461 0.7764 1.0
Qwen2.5-7B, mT0-Large, FLAN-T5-Large 0.7841 0.8165 0.8162 0.0
@@
Ensemble BLEU ROUGE-1 ROUGE-2 ROUGE-L BLEURT SBERT
mT0 Large, Qwen2.5-1.5B,
FLAN-T5 Large
0.8235 0.8241 0.8128 0.8247 0.8193 0.8235
mT0 Large, Qwen2.5-1.5B,
FLAN-T5 Large, Llama-3.2-1B
0.8169 0.822 0.8146 0.823 0.8143 0.8144
Qwen2.5-1.5B, FLAN-T5 Large,
Gemma3-1B
0.7889 0.7938 0.778 0.7944 0.7885 0.789
Qwen2.5-1.5B, FLAN-T5 Large,
Gemma3-1B, Llama-3.2-1B
0.7809 0.7916 0.7757 0.7915 0.788 0.7814
Qwen2.5-1.5B, FLAN-T5-Large,
Llama3.2-1B
0.7899 0.7937 0.7762 0.7944 0.7889 0.792
Llama3.2-3B, Qwen2.5-3B,
FLAN-T5-Large
0.8115 0.8151 0.8057 0.8151 0.8168 0.812
Qwen2.5-7B, mT0-Large,
FLAN-T5-Large
0.8333 0.8325 0.8253 0.833 0.8334 0.8353
Qwen2.5-7B, mT0-Large,
Llama3.2-3B
0.8272 0.8314 0.8248 0.831 0.8299 0.8297
Qwen2.5-7B, Llama3.1-8B,
mT0-Large
0.8315 0.8312 0.8262 0.8302 0.8324 0.8321
Llama3.1-8B, mT0-Large,
Qwen2.5-1.5B
0.8259 0.8254 0.8192 0.8251 0.827 0.8262
@@ H.2.6 TriviaQA
Model Random Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile
Qwen2.5-7B 0.6575 0.7258 0.7238 0.7242 0.1
Qwen2.5-3B 0.6003 0.6684 0.6662 0.6664 0.1
Qwen2.5-1.5B 0.5556 0.6168 0.6147 0.6154 0.1
Qwen2.5-0.5B 0.4569 0.4832 0.4735 0.4812 0.0
Gemma3-1B 0.5063 0.5452 0.544 0.5442 0.0
Llama3.2-1B 0.5265 0.5745 0.5708 0.572 0.1
Llama3.2-3B 0.6584 0.7202 0.7183 0.7193 0.1
Llama3.1-8B 0.7303 0.7741 0.7711 0.7714 0.1
FLAN-T5-Large 0.4761 0.5092 0.5019 0.5102 0.0
mT0-Large 0.4337 0.4464 0.4436 0.4479 0.2
@@
Ensemble Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile
Qwen2.5-1.5B, Llama3.2-1B 0.612 0.6183 0.6155 0.1
Llama3.2-1B, Gemma3-1B 0.5763 0.5788 0.5782 0.1
Qwen2.5-1.5B, Llama3.2-1B, Gemma3-1B 0.6079 0.6114 0.6111 0.1
Qwen2.5-1.5B, Llama3.2-1B, FLAN-T5-Large 0.5917 0.6002 0.6 0.0
Llama3.1-8B, Qwen2.5-7B 0.7596 0.7534 0.763 0.0
Llama3.1-8B, Qwen2.5-7B, mT0-Large 0.6443 0.6188 0.6618 0.0
Qwen2.5-7B, Llama3.2-3B 0.7402 0.7434 0.7421 0.1
Qwen2.5-7B, Llama3.2-3B, Gemma3-1B 0.689 0.6737 0.6939 0.0
Qwen2.5-7B, Gemma3-1B 0.6433 0.6429 0.6493 0.0
Llama3.2-3B, Qwen2.5-3B 0.7128 0.7119 0.7144 0.1
Llama3.1-8B, Llama3.2-3B, Qwen2.5-3B 0.7398 0.731 0.7455 0.1
@@
Ensemble BLEU ROUGE-1 ROUGE-2 ROUGE-L BLEURT SBERT
Qwen2.5-1.5B, Llama3.2-1B,
Gemma3-1B
0.6 0.6054 0.5699 0.6056 0.5966 0.5934
Qwen2.5-1.5B, Llama3.2-1B,
FLAN-T5-Large
0.5897 0.5963 0.565 0.5965 0.5849 0.5838
Llama3.2-1B, Gemma3-1B,
FLAN-T5-Large
0.5589 0.5656 0.5342 0.5652 0.5585 0.5551
Qwen2.5-1.5B, Llama3.2-1B,
Gemma3-1B, FLAN-T5-Large
0.5984 0.6027 0.5676 0.6026 0.5963 0.5885
Llama3.1-8B, Qwen2.5-7B,
Llama3.2-3B
0.7684 0.7718 0.744 0.772 0.7653 0.7584
Qwen2.5-7B, Llama3.2-3B,
Gemma3-1B
0.7169 0.7211 0.6823 0.7216 0.7146 0.7078
Qwen2.5-7B, Llama3.2-3B,
Qwen2.5-3B
0.729 0.7343 0.6881 0.7342 0.7322 0.7259
Llama3.1-8B, Llama3.2-3B,
Qwen2.5-3B
0.7622 0.7653 0.7422 0.7653 0.7551 0.7482
Llama3.1-8B, Qwen2.5-7B,
mT0-Large
0.7525 0.755 0.7408 0.7551 0.7319 0.7194
@@ H.3 Latencies at 98% Target Model Performance
Model Random Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile
Qwen2.5-7B 3136 2495 2271 2310 0.1
Qwen2.5-3B 2725 2384 2258 2244 0.3
Qwen2.5-1.5B 2541 2268 2093 2127 0.2
Qwen2.5-0.5B 2416 2331 2300 2300 0.2
Gemma3-1B 2480 2233 2147 2126 0.2
Llama3.2-1B 2451 2316 2296 2283 0.0
Llama3.2-3B 2680 2313 2134 2151 0.2
Llama3.1-8B 3177 2278 2097 2055 0.2
FLAN-T5-Large 2469 2340 2340 2339 0.0
mT0-Large $26$ 2484 2443 2449 0.3
@@
Ensemble Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile
Gemma3-1B, Qwen2.5-1.5B 2256 2102 2125 0.1
Gemma3-1B, Qwen2.5-1.5B, Llama3.2-1B 2239 2217 2220 0.1
Gemma3-1B, Qwen2.5-1.5B, mT0-Large 2370 2290 2300 0.1
Gemma3-1B, Qwen2.5-1.5B, FLAN-T5-Large 2273 2259 2236 0.1
Qwen2.5-1.5B, Llama3.2-1B, mT0-Large 2395 2386 2383 0.2
Llama3.1-8B, Qwen2.5-7B 2246 2117 2129 0.2
Llama3.1-8B, Qwen2.5-3B 2620 2388 2362 0.6
Llama3.1-8B, Qwen2.5-7B, FLAN-T5-Large 2745 2773 2761 0.1
Llama3.1-8B, Qwen2.5-7B, mT0-Large 2773 2696 2376 1.0
@@
Ensemble BLEU ROUGE-1 ROUGE-2 ROUGE-L BLEURT SBERT
Gemma3-1B, Qwen2.5-1.5B,
Llama3.2-1B
2223 2239 2245 2239 2001 2215
Gemma3-1B, Qwen2.5-1.5B,
mT0-Large
2286 2296 2293 2280 2077 2282
Qwen2.5-1.5B, Gemma3-1B,
FLAN-T5-Large
2223 2237 2197 2187 1931 2156
Qwen2.5-1.5B, Llama3.2-1B,
mT0-Large
2359 2342 2361 2350 2162 2324
Gemma3-1B, Qwen2.5-1.5B,
Llama3.2-1B, mT0-Large
2272 2308 2293 2293 2033 2257
Gemma3-1B, Qwen2.5-1.5B,
mT0-Large, FLAN-T5-Large
2294 2276 2279 2251 1996 2238
Llama3.1-8B, Qwen2.5-7B,
FLAN-T5-Large
2336 2343 2233 2297 1737 2252
Llama3.1-8B, Qwen2.5-7B,
mT0-Large
2336 2315 2376 2329 1557 2301
Llama3.1-8B, Qwen2.5-7B,
Qwen2.5-1.5B
2311 2286 2275 2252 1693 2171
Llama3.2-3B, Qwen2.5-3B,
Gemma3-1B
2218 2225 2211 2188 1856 2131
@@
Model Random Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile
Qwen2.5-7B 1767 499.4 499.4 499.4 0.0
Qwen2.5-3B 1530 394 318.3 313 1.0
Qwen2.5-1.5B 1428 436.9 324.2 335.7 0.1
Qwen2.5-0.5B 1346 951 885.9 906.1 0.2
Gemma3-1B 1396 811 716.8 722.1 0.5
Llama3.2-1B 1373 1076 1077 1014 0.0
Llama3.2-3B 1505 357.4 309.9 283.5 0.6
Llama3.1-8B 1784 516.1 516.1 516.1 0.0
FLAN-T5-Large 1370 943.5 787.7 806.2 0.2
mT0-Large 1495 1208 $11$ 1130 0.1
@@
Ensemble Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile
Gemma3-1B, Qwen2.5-1.5B 597.1 575.1 416.7 0.8
Gemma3-1B, Qwen2.5-1.5B, Llama3.2-1B 809.2 866.5 835.7 0.0
Gemma3-1B, Qwen2.5-1.5B, mT0-Large 890 852.2 826.7 0.2
Gemma3-1B, Qwen2.5-1.5B, FLAN-T5-Large 643.8 636.7 653.4 0.3
Qwen2.5-1.5B, Llama3.2-1B, mT0-Large 916.4 988.6 986 0.0
Llama3.1-8B, Qwen2.5-7B 516.1 516.1 516.1 0.0
Llama3.1-8B, Qwen2.5-3B 516.1 516.1 516.1 0.0
Llama3.1-8B, Qwen2.5-7B, FLAN-T5-Large 516.1 565.4 516.1 0.2
Llama3.1-8B, Qwen2.5-7B, mT0-Large 724.7 883.2 516.1 1.0
@@
Ensemble BLEU ROUGE-1 ROUGE-2 ROUGE-L BLEURT SBERT
Gemma3-1B, Qwen2.5-1.5B,
Llama3.2-1B
575.1 552.2 556.6 546.9 309.3 519.6
Gemma3-1B, Qwen2.5-1.5B,
mT0-Large
590.8 527.4 617.2 503.6 324.9 473.7
Qwen2.5-1.5B, Gemma3-1B,
FLAN-T5-Large
437.8 414 443.1 399 194.8 358.6
Qwen2.5-1.5B, Llama3.2-1B,
mT0-Large
737.8 644.4 725.4 678.8 479 647.1
Gemma3-1B, Qwen2.5-1.5B,
Llama3.2-1B, mT0-Large
610.1 574.9 622.4 551.1 324.9 572.3
Gemma3-1B, Qwen2.5-1.5B,
mT0-Large, FLAN-T5-Large
575.8 511.5 596.9 519.5 238.7 466.6
Llama3.1-8B, Qwen2.5-7B,
FLAN-T5-Large
516.1 516.1 516.1 516.1 516.1 516.1
Llama3.1-8B, Qwen2.5-7B,
mT0-Large
516.1 516.1 516.1 516.1 516.1 516.1
Llama3.1-8B, Qwen2.5-7B,
Qwen2.5-1.5B
516.1 516.1 516.1 516.1 516.1 516.1
Llama3.2-3B, Qwen2.5-3B,
Gemma3-1B
290.1 345.5 326.2 319.1 262.8 262.8
@@
Model Random Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile
Qwen2.5-7B 2583 1672 1460 1556 0.1
Qwen2.5-3B 2229 1817 1659 1696 0.1
Qwen2.5-1.5B 2083 1690 1547 1590 0.1
Qwen2.5-0.5B 1979 1845 1797 1794 0.5
Gemma3-1B 2051 1630 1489 1536 0.1
Llama3.2-1B 2007 1673 1581 1584 0.2
Llama3.2-3B 2205 1578 1347 1356 0.1
Llama3.1-8B 2604 1528 1215 1166 0.6
FLAN-T5-Large 2037 1902 1897 1864 0.1
mT0-Large 2170 1959 1912 1918 0.4

### v3

@@ H.2.4 CNN/DailyMail
Model Random Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile Qwen2.5-7B 0.258 0.2577 0.2605 0.2607 0.2 Qwen2.5-3B 0.2511 0.2539 0.2537 0.2537 0.5 Qwen2.5-1.5B 0.2474 0.2407 0.2536 0.2559 1.0 Qwen2.5-0.5B 0.2418 0.2371 0.245 0.2467 1.0 Gemma3-1B 0.2448 0.2454 0.2473 0.2479 0.4 Llama3.2-1B 0.2482 0.2502 0.2508 0.2511 0.3 Llama3.2-3B 0.2522 0.2535 0.2548 0.2549 0.1 Llama3.1-8B 0.2551 0.2547 0.2564 0.2565 0.2 mT0-Large 0.2527 0.2528 0.2539 0.2542 0.8
@@
Ensemble Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile Llama3.2-1B, mT0-Large 0.2554 0.2529 0.2553 0.1 Llama3.2-1B, Qwen2.5-1.5B 0.2532 0.2493 0.2548 0.2 Qwen2.5-1.5B, Gemma3-1B 0.2505 0.2489 0.2559 1.0 Llama3.2-1B, mT0-Large, Qwen2.5-1.5B 0.2568 0.2491 0.257 0.8 Qwen2.5-7B, Llama3.1-8B 0.2579 0.259 0.2611 0.2 Qwen2.5-7B, mT0-Large 0.258 0.2575 0.2588 0.2 Qwen2.5-7B, Qwen2.5-3B 0.259 0.2583 0.2602 0.2
@@
Ensemble BLEU ROUGE-1 ROUGE-2 ROUGE-L BLEURT SBERT Llama3.2-1B, mT0-Large, Qwen2.5-1.5B 0.2611 0.2616 0.2598 0.2608 0.2567 0.2603 Llama3.2-1B, mT0-Large, Qwen2.5-0.5B 0.2566 0.2567 0.2553 0.2574 0.2513 0.2547 Qwen2.5-1.5B, Qwen2.5-0.5B, Gemma3-1B 0.2551 0.2546 0.2514 0.2519 0.2522 0.2515 Llama3.2-1B, Gemma3-1B, Qwen2.5-0.5B 0.2537 0.2538 0.2534 0.2543 0.251 0.2528 Llama3.2-3B, Qwen2.5-3B, Qwen2.5-0.5B 0.2568 0.258 0.2578 0.2578 0.2546 0.2576 Qwen2.5-7B, mT0-Large, Llama3.2-3B 0.2635 0.2618 0.2615 0.2607 0.2584 0.2618 Qwen2.5-7B, Llama3.1-8B, mT0-Large 0.2622 0.2613 0.2622 0.2619 0.2586 0.2612 Llama3.1-8B, mT0-Large, Qwen2.5-3B 0.2589 0.2571 0.2583 0.2578 0.2555 0.2585
@@
Model Random Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile Qwen2.5-7B 0.8053 0.8217 0.8201 0.821 0.0 Qwen2.5-3B 0.7623 0.7916 0.7921 0.7915 0.1 Qwen2.5-1.5B 0.7593 0.7897 0.7847 0.7868 0.0 Qwen2.5-0.5B 0.642 0.6898 0.6826 0.685 0.0 Llama3.1-8B 0.7987 0.8173 0.8151 0.8169 0.0 Llama3.2-3B 0.7627 0.7764 0.7791 0.7833 0.5 Llama3.2-1B 0.631 0.6695 0.677 0.6768 0.3 mT0-Large 0.802 0.8205 0.8211 0.8211 0.1 FLAN-T5-Large 0.7123 0.7598 0.6896 0.7437 0.0 Gemma3-1B 0.676 0.6984 0.6938 0.696 0.0
@@
Ensemble Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile mT0-Large, Qwen2.5-1.5B 0.8155 0.8168 0.8125 0.0 mT0-Large, Qwen2.5-1.5B, FLAN-T5-Large 0.7809 0.8081 0.8043 0.0 Qwen2.5-1.5B, FLAN-T5-Large, Gemma3-1B 0.7414 0.7511 0.7632 0.0 Qwen2.5-7B, Llama3.1-8B 0.8211 0.8211 0.8219 0.0 Qwen2.5-7B, Llama3.1-8B, mT0-Large 0.8236 0.8239 0.8255 0.0 Qwen2.5-7B, Llama3.2-3B 0.7993 0.7985 0.7988 0.0 Qwen2.5-7B, Llama3.2-3B, Gemma3-1B 0.7599 0.7461 0.7764 1.0 Qwen2.5-7B, mT0-Large, FLAN-T5-Large 0.7841 0.8165 0.8162 0.0
@@
Ensemble BLEU ROUGE-1 ROUGE-2 ROUGE-L BLEURT SBERT mT0 Large, Qwen2.5-1.5B, FLAN-T5 Large 0.8235 0.8241 0.8128 0.8247 0.8193 0.8235 mT0 Large, Qwen2.5-1.5B, FLAN-T5 Large, Llama-3.2-1B 0.8169 0.822 0.8146 0.823 0.8143 0.8144 Qwen2.5-1.5B, FLAN-T5 Large, Gemma3-1B 0.7889 0.7938 0.778 0.7944 0.7885 0.789 Qwen2.5-1.5B, FLAN-T5 Large, Gemma3-1B, Llama-3.2-1B 0.7809 0.7916 0.7757 0.7915 0.788 0.7814 Qwen2.5-1.5B, FLAN-T5-Large, Llama3.2-1B 0.7899 0.7937 0.7762 0.7944 0.7889 0.792 Llama3.2-3B, Qwen2.5-3B, FLAN-T5-Large 0.8115 0.8151 0.8057 0.8151 0.8168 0.812 Qwen2.5-7B, mT0-Large, FLAN-T5-Large 0.8333 0.8325 0.8253 0.833 0.8334 0.8353 Qwen2.5-7B, mT0-Large, Llama3.2-3B 0.8272 0.8314 0.8248 0.831 0.8299 0.8297 Qwen2.5-7B, Llama3.1-8B, mT0-Large 0.8315 0.8312 0.8262 0.8302 0.8324 0.8321 Llama3.1-8B, mT0-Large, Qwen2.5-1.5B 0.8259 0.8254 0.8192 0.8251 0.827 0.8262
@@ H.2.6 TriviaQA
Model Random Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile Qwen2.5-7B 0.6575 0.7258 0.7238 0.7242 0.1 Qwen2.5-3B 0.6003 0.6684 0.6662 0.6664 0.1 Qwen2.5-1.5B 0.5556 0.6168 0.6147 0.6154 0.1 Qwen2.5-0.5B 0.4569 0.4832 0.4735 0.4812 0.0 Gemma3-1B 0.5063 0.5452 0.544 0.5442 0.0 Llama3.2-1B 0.5265 0.5745 0.5708 0.572 0.1 Llama3.2-3B 0.6584 0.7202 0.7183 0.7193 0.1 Llama3.1-8B 0.7303 0.7741 0.7711 0.7714 0.1 FLAN-T5-Large 0.4761 0.5092 0.5019 0.5102 0.0 mT0-Large 0.4337 0.4464 0.4436 0.4479 0.2
@@
Ensemble Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile Qwen2.5-1.5B, Llama3.2-1B 0.612 0.6183 0.6155 0.1 Llama3.2-1B, Gemma3-1B 0.5763 0.5788 0.5782 0.1 Qwen2.5-1.5B, Llama3.2-1B, Gemma3-1B 0.6079 0.6114 0.6111 0.1 Qwen2.5-1.5B, Llama3.2-1B, FLAN-T5-Large 0.5917 0.6002 0.6 0.0 Llama3.1-8B, Qwen2.5-7B 0.7596 0.7534 0.763 0.0 Llama3.1-8B, Qwen2.5-7B, mT0-Large 0.6443 0.6188 0.6618 0.0 Qwen2.5-7B, Llama3.2-3B 0.7402 0.7434 0.7421 0.1 Qwen2.5-7B, Llama3.2-3B, Gemma3-1B 0.689 0.6737 0.6939 0.0 Qwen2.5-7B, Gemma3-1B 0.6433 0.6429 0.6493 0.0 Llama3.2-3B, Qwen2.5-3B 0.7128 0.7119 0.7144 0.1 Llama3.1-8B, Llama3.2-3B, Qwen2.5-3B 0.7398 0.731 0.7455 0.1
@@
Ensemble BLEU ROUGE-1 ROUGE-2 ROUGE-L BLEURT SBERT Qwen2.5-1.5B, Llama3.2-1B, Gemma3-1B 0.6 0.6054 0.5699 0.6056 0.5966 0.5934 Qwen2.5-1.5B, Llama3.2-1B, FLAN-T5-Large 0.5897 0.5963 0.565 0.5965 0.5849 0.5838 Llama3.2-1B, Gemma3-1B, FLAN-T5-Large 0.5589 0.5656 0.5342 0.5652 0.5585 0.5551 Qwen2.5-1.5B, Llama3.2-1B, Gemma3-1B, FLAN-T5-Large 0.5984 0.6027 0.5676 0.6026 0.5963 0.5885 Llama3.1-8B, Qwen2.5-7B, Llama3.2-3B 0.7684 0.7718 0.744 0.772 0.7653 0.7584 Qwen2.5-7B, Llama3.2-3B, Gemma3-1B 0.7169 0.7211 0.6823 0.7216 0.7146 0.7078 Qwen2.5-7B, Llama3.2-3B, Qwen2.5-3B 0.729 0.7343 0.6881 0.7342 0.7322 0.7259 Llama3.1-8B, Llama3.2-3B, Qwen2.5-3B 0.7622 0.7653 0.7422 0.7653 0.7551 0.7482 Llama3.1-8B, Qwen2.5-7B, mT0-Large 0.7525 0.755 0.7408 0.7551 0.7319 0.7194
@@ H.3 Latencies at 98% Target Model Performance
Model Random Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile Qwen2.5-7B 3136 2495 2271 2310 0.1 Qwen2.5-3B 2725 2384 2258 2244 0.3 Qwen2.5-1.5B 2541 2268 2093 2127 0.2 Qwen2.5-0.5B 2416 2331 2300 2300 0.2 Gemma3-1B 2480 2233 2147 2126 0.2 Llama3.2-1B 2451 2316 2296 2283 0.0 Llama3.2-3B 2680 2313 2134 2151 0.2 Llama3.1-8B 3177 2278 2097 2055 0.2 FLAN-T5-Large 2469 2340 2340 2339 0.0 mT0-Large 2626 2484 2443 2449 0.3
@@
Ensemble Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile Gemma3-1B, Qwen2.5-1.5B 2256 2102 2125 0.1 Gemma3-1B, Qwen2.5-1.5B, Llama3.2-1B 2239 2217 2220 0.1 Gemma3-1B, Qwen2.5-1.5B, mT0-Large 2370 2290 2300 0.1 Gemma3-1B, Qwen2.5-1.5B, FLAN-T5-Large 2273 2259 2236 0.1 Qwen2.5-1.5B, Llama3.2-1B, mT0-Large 2395 2386 2383 0.2 Llama3.1-8B, Qwen2.5-7B 2246 2117 2129 0.2 Llama3.1-8B, Qwen2.5-3B 2620 2388 2362 0.6 Llama3.1-8B, Qwen2.5-7B, FLAN-T5-Large 2745 2773 2761 0.1 Llama3.1-8B, Qwen2.5-7B, mT0-Large 2773 2696 2376 1.0
@@
Ensemble BLEU ROUGE-1 ROUGE-2 ROUGE-L BLEURT SBERT Gemma3-1B, Qwen2.5-1.5B, Llama3.2-1B 2223 2239 2245 2239 2001 2215 Gemma3-1B, Qwen2.5-1.5B, mT0-Large 2286 2296 2293 2280 2077 2282 Qwen2.5-1.5B, Gemma3-1B, FLAN-T5-Large 2223 2237 2197 2187 1931 2156 Qwen2.5-1.5B, Llama3.2-1B, mT0-Large 2359 2342 2361 2350 2162 2324 Gemma3-1B, Qwen2.5-1.5B, Llama3.2-1B, mT0-Large 2272 2308 2293 2293 2033 2257 Gemma3-1B, Qwen2.5-1.5B, mT0-Large, FLAN-T5-Large 2294 2276 2279 2251 1996 2238 Llama3.1-8B, Qwen2.5-7B, FLAN-T5-Large 2336 2343 2233 2297 1737 2252 Llama3.1-8B, Qwen2.5-7B, mT0-Large 2336 2315 2376 2329 1557 2301 Llama3.1-8B, Qwen2.5-7B, Qwen2.5-1.5B 2311 2286 2275 2252 1693 2171 Llama3.2-3B, Qwen2.5-3B, Gemma3-1B 2218 2225 2211 2188 1856 2131
@@
Model Random Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile Qwen2.5-7B 1767 499.4 499.4 499.4 0.0 Qwen2.5-3B 1530 394 318.3 313 1.0 Qwen2.5-1.5B 1428 436.9 324.2 335.7 0.1 Qwen2.5-0.5B 1346 951 885.9 906.1 0.2 Gemma3-1B 1396 811 716.8 722.1 0.5 Llama3.2-1B 1373 1076 1077 1014 0.0 Llama3.2-3B 1505 357.4 309.9 283.5 0.6 Llama3.1-8B 1784 516.1 516.1 516.1 0.0 FLAN-T5-Large 1370 943.5 787.7 806.2 0.2 mT0-Large 1495 1208 1111 1130 0.1
@@
Ensemble Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile Gemma3-1B, Qwen2.5-1.5B 597.1 575.1 416.7 0.8 Gemma3-1B, Qwen2.5-1.5B, Llama3.2-1B 809.2 866.5 835.7 0.0 Gemma3-1B, Qwen2.5-1.5B, mT0-Large 890 852.2 826.7 0.2 Gemma3-1B, Qwen2.5-1.5B, FLAN-T5-Large 643.8 636.7 653.4 0.3 Qwen2.5-1.5B, Llama3.2-1B, mT0-Large 916.4 988.6 986 0.0 Llama3.1-8B, Qwen2.5-7B 516.1 516.1 516.1 0.0 Llama3.1-8B, Qwen2.5-3B 516.1 516.1 516.1 0.0 Llama3.1-8B, Qwen2.5-7B, FLAN-T5-Large 516.1 565.4 516.1 0.2 Llama3.1-8B, Qwen2.5-7B, mT0-Large 724.7 883.2 516.1 1.0
@@
Ensemble BLEU ROUGE-1 ROUGE-2 ROUGE-L BLEURT SBERT Gemma3-1B, Qwen2.5-1.5B, Llama3.2-1B 575.1 552.2 556.6 546.9 309.3 519.6 Gemma3-1B, Qwen2.5-1.5B, mT0-Large 590.8 527.4 617.2 503.6 324.9 473.7 Qwen2.5-1.5B, Gemma3-1B, FLAN-T5-Large 437.8 414 443.1 399 194.8 358.6 Qwen2.5-1.5B, Llama3.2-1B, mT0-Large 737.8 644.4 725.4 678.8 479 647.1 Gemma3-1B, Qwen2.5-1.5B, Llama3.2-1B, mT0-Large 610.1 574.9 622.4 551.1 324.9 572.3 Gemma3-1B, Qwen2.5-1.5B, mT0-Large, FLAN-T5-Large 575.8 511.5 596.9 519.5 238.7 466.6 Llama3.1-8B, Qwen2.5-7B, FLAN-T5-Large 516.1 516.1 516.1 516.1 516.1 516.1 Llama3.1-8B, Qwen2.5-7B, mT0-Large 516.1 516.1 516.1 516.1 516.1 516.1 Llama3.1-8B, Qwen2.5-7B, Qwen2.5-1.5B 516.1 516.1 516.1 516.1 516.1 516.1 Llama3.2-3B, Qwen2.5-3B, Gemma3-1B 290.1 345.5 326.2 319.1 262.8 262.8
@@
Model Random Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile Qwen2.5-7B 2583 1672 1460 1556 0.1 Qwen2.5-3B 2229 1817 1659 1696 0.1 Qwen2.5-1.5B 2083 1690 1547 1590 0.1 Qwen2.5-0.5B 1979 1845 1797 1794 0.5 Gemma3-1B 2051 1630 1489 1536 0.1 Llama3.2-1B 2007 1673 1581 1584 0.2 Llama3.2-3B 2205 1578 1347 1356 0.1 Llama3.1-8B 2604 1528 1215 1166 0.6 FLAN-T5-Large 2037 1902 1897 1864 0.1 mT0-Large 2170 1959 1912 1918 0.4